OpenStudy (anonymous):

solve 3(x+4)-4(3x-10)=4(2x-3)+6(3x+1

OpenStudy (anonymous):

Solve LHS and RHS. And you'll get the result.

OpenStudy (anonymous):

@traceyt: \[3(x+4) -4(3x-10) = 4(2x+3) + 6(3x +1)\] \[= 3x + 12 - 12x + 40 = 8x + 12 +18x + 6\] \[-9x +52 = 26x + 18\] \[34x = 34 => x = \frac{ 34 }{ 35}\]